# AFABench: A Generic Framework for Benchmarking Active Feature Acquisition

### v1

@@ 3.1. Methods
Table 1. Summary of feature selection methods included in the benchmark. The possible categories include generative greedy ($G$), discriminative greedy (DG), model-free RL (MFRL), model-based RL (MBRL), and static selection (S).
Paradigm Strategy Greedy? Representative Method(s)
AFA (Dynamic Feature Selection) Generative estimation of CMI Greedy EDDI-GG (Ma et al., 2019)
Discriminative estimation of CMI Greedy GDFS-DG (Covert et al., 2023)
DIME-DG (Gadgil et al., 2024)
Model-free RL Non-greedy JAFA-MFRL (Shim et al., 2018)
@@
Model-based RL Non-greedy ODIN-MBRL (Zannone et al., 2019)
Oracle-based Non-Greedy AACO (Valancius et al., 2024)
@@
As described in the recent survey by Rahbar et al., AFA methods can be broadly categorized into two main types: (i) greedy approaches, which iteratively select the feature that provides the highest expected information gain given the features already acquired. These greedy methods can be further classified as discriminative or generative; and (ii) non-greedy approaches, which aim to learn feature selection policies typically using reinforcement learning (RL) techniques. Non-greedy methods can be further divided into model-free RL and model-based RL approaches.
@@
In this benchmark, we have selected at least one representative state-of-the-art method from each category to provide a comprehensive evaluation of the various strategies for AFA. Table 3.1 summarizes the methods included in the benchmark and their corresponding categories. In addition, we include two static feature selection methods to highlight the potential benefits of dynamic selection. It should be noted that some existing methods do not fit neatly into any of the above categories (Rahbar et al., 2025). We discuss these methods and the rationale for excluding them from this benchmark in Section 3.4. In the following sections, we briefly describe each of the included methods.
3.1.1. Greedy methods
To quantify the information gain of a candidate feature conditioned on those already selected (as done in greedy methods), conditional mutual information (CMI) is commonly used. CMI measures the mutual dependence between a feature and the target variable, given the features observed so far. Formally, the optimal selection policy is one that chooses the feature $\mathbf{x}_{i}$ that maximizes the mutual information with the target $\mathbf{y}$, conditioned on the current set of observed features $x_{S}$. This quantity, denoted as $I(\mathbf{y};\mathbf{x}_{i}\mid x_{S})$, is defined using the Kullback–Leibler divergence (Cover and Thomas, 2012; Covert et al., 2023):
(2) $I(\mathbf{y};\mathbf{x}_{i}\mid x_{S})=D_{\text{KL}}\big{(}p(\mathbf{x}_{i},\mathbf{y}\mid x_{S})\,\big{\|}\,p(\mathbf{x}_{i}\mid x_{S})\,p(\mathbf{y}\mid x_{S})\big{)}.$
The corresponding feature selection policy is then given by:
(3) $\pi(x_{S})=\operatorname*{argmax}_{i\in[n]}I(\mathbf{y};\mathbf{x}_{i}\mid x_{S}).$
@@ 3.1.1. Greedy methods
In practice, implementing this ideal policy is challenging. At each decision point, computing $I(\mathbf{y};\mathbf{x}_{i}\mid x_{S})$ for all feature indices $i$ requires access to the joint and conditional distributions of the features and the target, conditioned on the observed subset. In the following, we describe how the included greedy methods approximate or sidestep this computation.
@@
EDDI-GG (Ma et al., 2019). EDDI estimates the CMI acquisition policy using a generative approach. At its core lies a novel partial Variational Autoencoder (PVAE) capable of handling arbitrarily missing features. Once trained on the full dataset $\mathcal{D}$, the PVAE offers fast approximations of posterior predictive distributions such as $p(\mathbf{x}_{i}\mid x_{S})$: it samples a latent code conditioned on the observed subset $x_{S}$ and subsequently generates the candidate feature $\mathbf{x}_{i}$. Whenever these predictive distributions admit a Gaussian parameterization, the CMI can be computed analytically; otherwise, it is efficiently estimated via Monte Carlo sampling. A limitation of this approach is that the PVAE is expensive to train. This limitation is addressed in the next two discriminative approaches.
@@
GDFS-DG (Covert et al., 2023). This approach formulates an optimization problem that evaluates a policy by its one-step-ahead prediction accuracy. Specifically, when the policy has chosen the next feature $i$ given the currently observed set $x_{S}$, the goal is to minimize the expected loss $\mathbb{E}\!\left[\,l\bigl{(}f(x_{S},\mathbf{x}_{i}),\mathbf{y}\bigr{)}\right]$, where $l$ measures the discrepancy between the prediction and the target $\mathbf{y}$. The optimal predictor for this objective is the Bayes classifier, $f^{*}(x_{S})=p(\mathbf{y}\mid x_{S})$. Moreover, under this classifier, the policy that minimizes the expected one-step-ahead loss coincides with the CMI policy of Eq. (3). The resulting optimization is solved with amortized optimization (Amos, 2022): the variational objective is expressed in a closed form and then minimized using a deep neural network.
@@
DIME-DG (Gadgil et al., 2024). This approach directly estimates the CMI in a discriminative way. Extending the variational estimators of Chattopadhyay et al.; Covert et al., introduces a new objective whose optimum equals the true CMI. The practical implementation uses two jointly trained networks: a predictor $f_{\theta}$ and a value network $g_{\phi}$ that outputs $I(\mathbf{y};\mathbf{x}_{i}\mid x_{S})$ for each unobserved feature. When $f_{\theta}$ approaches the Bayes classifier, the expected reduction in cross-entropy achieved by adding the feature $i$ matches its CMI; training $g_{\phi}$ to reproduce this reduction yields a consistent estimator. After training, $g_{\phi}$ selects the next feature by choosing the largest predicted CMI.
3.1.2. Non-greedy methods via reinforcement learning
Given the sequential nature of active feature acquisition, framing it as a Markov Decision Process (MDP) offers a principled foundation. In this setting, reinforcement learning can be used to train non-greedy policies $\pi$ that go beyond immediate information gain, selecting features that maximize long-term reward. This approach enables the discovery of acquisition strategies that can outperform myopic and greedy methods by considering the cumulative benefit of future selections. We begin by describing the standard MDP formulation of the AFA problem for the hard-budget case (Janisch et al., 2020).
(4) $s$ $=s(S,x_{S})$
$a$ $\in U,\quad U=\{1,\dots,n\}\setminus S$
$r(s,a)$ $=r(f(x_{S}),f(x_{S\cup a}),y,\mathbb{I}(|S|=b))$
Here, $S$ are the indices of the features observed currently and $x_{S}$ are the features observed. The state $s$ is generally a function of these two vectors. In some cases, it is a simple concatenation of the two vectors, but more sophisticated methods often use special neural network architectures for encoding sets.
The discrete action space $U$ comprises the set of feature indices not yet selected. Note that we exclude the stop action often present in other papers since we focus solely on the hard-budget case.
The reward received after choosing action $a$ in state $s$, $r(s,a)$, is generally a function of the current classifier predictions, the next classifier predictions, the true label and whether the episode is finished or not. In many cases, the negative classification loss $l(f(x_{S\cup a},y)$ is used.
If one uses the reward $r(s,a)=-l(f(x_{S\cup a},y)$ without discounting, then the resulting MDP corresponds directly to the optimization objective in Eq. (1), under two simplifying assumptions: (i) a fixed budget $b$ on the number of features and (ii) uniform acquisition cost across all features (see Section 2).
@@ 3.1.2. Non-greedy methods via reinforcement learning
JAFA-MFRL (Shim et al., 2018). This method uses the RNN-based and order-invariant set encoding introduced in Vinyals et al. to represent the subsets of acquired features, significantly reducing the complexity of the search space. The approach jointly trains a reinforcement learning agent and a classifier, ensuring that the feature selection policy is optimized with respect to the classification objective. To solve the underlying MDP, the method uses deep Q-Networks (DQN) (Mnih et al., 2015), enabling efficient policy learning even in high-dimensional feature spaces.
Adopted for the hard-budget case and with uniform acquisition costs, the reward function for this method is the negative classification loss, but only applied at the end of the episode:
(5) $r(s,a)=-l\big{(}f(x_{S\cup a}),y\big{)}\,\mathbb{I}(|S\cup a|=b).$
This sparse reward does not incentivize the agent to select features in any specific order.
In particular, this is an model-free RL approach, in contrast to the next method.
@@
ODIN-MBRL (Zannone et al., 2019). This is a model-based RL framework for AFA. ODIN uses a pre-trained PVAE (Ma et al., 2019) to approximate the conditional distribution of unobserved features, $p(\mathbf{x}_{i}\mid x_{S})$, for $i\notin S$. Since MDP transition dynamics can be derived from this distribution, ODIN performs model-based rollouts to simulate feature acquisition trajectories, greatly improving data efficiency and robustness to missing features. The policy is trained using Proximal Policy Optimization (PPO) (Schulman et al., 2017).
@@
In contrast to JAFA-MFRL, this method uses the dense reward function
(6) $r(s,a)=-l\big{(}f(x_{S\cup a}),y\big{)}$
which motivates the agent to select the most informative features first.
@@
ODIN-MFRL The key strength of ODIN-MBRL is its performance on smaller datasets (Zannone et al., 2019). However, since most of the datasets in this benchmark are relatively large, we decided to also evaluate a variant of ODIN where the agent is trained directly on the dataset, without using generative rollouts.
@@
OL-MFRL (Kachuee et al., 2019). This is another model-free and DQN-based method. The main characteristic of it is the reward function
(7) $r(s,a)=\|\textit{Cert}(x_{S\cup a})-\textit{Cert}(x_{S})\|$
@@
which does not depend on the true label but only on the classifier predictions. Here Cert is a vector of class probabilities, calculated by averaging over multiple forward passes with dropout.
Furthermore, the method uses a coupled PQ-network where the output of the predictor (P-network) is fed into the action-value estimator (Q-network).
3.1.3. Oracle-based methods
@@
Rather than learning policies through reinforcement learning, oracle-based methods directly approximate an optimal acquisition policy by estimating the underlying data distributions. This approach offers a middle ground between greedy methods and computationally expensive RL training.
@@
AACO (Valancius et al., 2024). The Acquisition Conditioned Oracle (ACO) method formulates AFA as a subset optimization problem. Instead of greedily selecting features one by one, ACO considers acquiring entire subsets of features simultaneously. The optimal policy selects the subset $o^{\prime}\subseteq\mathcal{O}$ that minimizes the expected acquisition cost plus the prediction loss,
@@
(8) $u(x_{o},o)=\operatorname*{argmin}_{o^{\prime}\subseteq\mathcal{O}}\mathbb{E}_{y,x_{o^{\prime}}\mid x_{o}}\left[\ell\left(\hat{y}(x_{o},x_{o^{\prime}}),y\right)\right]+\alpha|o^{\prime}|,$
@@
In practice, computing this expectation requires knowledge of the joint distribution $p(y,x_{o^{\prime}}\mid x_{o})$, which is unknown. The Acquisition Conditioned Approximate Oracle (AACO) addresses this by using $k$-nearest neighbors density estimation to approximate the required distributions. Furthermore, for computational tractability in high-dimensional spaces, AACO samples a random subset of possible feature combinations rather than evaluating all $2^{|\mathcal{O}|}$ possibilities.
This approach is non-greedy since it jointly optimizes over multiple features rather than selecting them sequentially, yet it avoids the training complexity of RL methods by directly approximating the optimal oracle policy.
3.1.4. Static feature selection methods

### v3

@@ Appendix D Detailed Description of Implemented AFA Methods
Table 2. Summary of feature selection methods included in the benchmark. The possible categories include generative myopic (GM), discriminative myopic (DM), model-free RL (MFRL), model-based RL (MBRL), and static selection (S).
Paradigm Strategy Myopic? Representative Method(s)
AFA (Dynamic Feature Selection) Generative estimation of CMI Myopic EDDI-GM (Ma et al., 2019)
Discriminative estimation of CMI Myopic GDFS-DM (Covert et al., 2023)
DIME-DM (Gadgil et al., 2024)
Model-free RL Non-myopic JAFA-MFRL (Shim et al., 2018)
@@
Model-based RL Non-myopic ODIN-MBRL (Zannone et al., 2019)
Oracle-based Non-myopic AACO (Valancius et al., 2024)
@@
As described in the recent survey by Aronsson et al. (2026), AFA methods can be broadly categorized as follows: (i) myopic heuristics, which perform no explicit planning beyond a one-step lookahead and are often further classified as discriminative or generative, (ii) non-myopic methods based on reinforcement learning (RL), which learn acquisition policies by exploiting the MDP structure underlying AFA (see Section˜2.4) and can be divided into model-free and model-based approaches, and (iii) non-myopic methods that do not rely on explicit RL.
@@
In this benchmark, we have selected at least one representative state-of-the-art method from each category to provide a comprehensive evaluation of the various strategies for AFA. Table 2 summarizes the methods included in the benchmark and their corresponding categories. In addition, we include two static feature selection methods to highlight the potential benefits of dynamic selection.
Below, we briefly describe each method and explain how it is adapted to the hard- and soft-budget settings. In the hard-budget setting, we use the same budget $B$ for all methods. In the soft-budget setting, we tune the threshold parameter $\alpha$ separately for each method, since it is implemented and interpreted slightly differently across methods (see the discussion below).
D.1. Myopic methods
@@
All myopic methods considered in this benchmark are based on the myopic CMI policy in (3) (i.e., $I(\mathbf{y};\mathbf{x}_{i}\mid x_{S})$). All methods in this category hence differ in how they compute this information gain. At each decision point, computing $I(\mathbf{y};\mathbf{x}_{i}\mid x_{S})$ for all feature indices $i$ requires access to the joint and conditional distributions of the features and the target, conditioned on the observed subset. In the following, we describe how the included myopic methods achieve this and how each are adapted to the hard/soft-budget setting.
For all myopic methods, unless stated otherwise, we adapt them to the hard- and soft-budget settings as follows. Under the hard-budget setting, the adaptation is immediate: we iteratively acquire features according to (3) until the total cost reaches the budget $B$. Under the soft-budget setting, we adopt a commonly used heuristic and continue acquiring features via (3) until it falls below a threshold for every remaining feature, that is, until $I(\mathbf{y};\mathbf{x}_{i}\mid\mathbf{x}_{S})/c_{i}<\alpha$ for all $i\in[d]\setminus S$ (Gadgil et al., 2024; Aronsson et al., 2026). To ensure consistency with the other methods in the benchmark, we slightly deviate from the original experimental setting for the discriminative myopic methods. For datasets with uniform costs, under the hard-budget setting, we use the same hard budget values for both training and evaluation, and under the soft-budget setting, we train these methods with the maximum training hard budget and then reuse them across different soft-budget thresholds. For datasets with non-uniform costs, we train the discriminative myopic methods with the maximum hard budget (i.e., all features) to avoid acquiring substantially more features during evaluation than during training.
@@
EDDI-GM (Ma et al., 2019). EDDI estimates the CMI acquisition policy using a generative approach. At its core lies a novel partial Variational Autoencoder (PVAE) capable of handling arbitrarily missing features. Once trained on the full dataset $\mathcal{D}$, the PVAE offers fast approximations of posterior predictive distributions such as $p(\mathbf{x}_{i}\mid x_{S})$: it samples a latent code conditioned on the observed subset $x_{S}$ and subsequently generates the candidate feature $\mathbf{x}_{i}$. Whenever these predictive distributions admit a Gaussian parameterization, the CMI can be computed analytically; otherwise, it is efficiently estimated via Monte Carlo sampling. A limitation of this approach is that the PVAE is expensive to train. This limitation is addressed in the next two discriminative approaches.
Adapting EDDI-GM to the hard-budget setting is straightforward, by simply acquiring features according to (3) until budget $B$ is reached. For the soft-budget setting, we use the following, commonly used heuristic: we acquire features according to $I(\mathbf{y};\mathbf{x}_{i}\mid x_{S})/x_{i}$ until $I(\mathbf{y};\mathbf{x}_{i}\mid x_{S})/x_{i}<\alpha$ for all $i\in[d]\setminus S$.
@@
GDFS-DM (Covert et al., 2023). This approach formulates an optimization problem that evaluates a policy by its one-step-ahead prediction accuracy. Specifically, when the policy has chosen the next feature $i$ given the currently observed set $x_{S}$, the goal is to minimize the expected loss $\mathbb{E}\left[\,l\bigl(f(x_{S},\mathbf{x}_{i}),\mathbf{y}\bigr)\right]$, where $l$ measures the discrepancy between the prediction and the target $\mathbf{y}$. The optimal predictor for this objective is the Bayes classifier, $f^{*}(x_{S})=p(\mathbf{y}\mid x_{S})$. Moreover, under this classifier, the policy that minimizes the expected one-step-ahead loss coincides with the CMI policy of Eq. (3). The resulting optimization is solved with amortized optimization (Amos, 2022): the variational objective is expressed in a closed form and then minimized using a deep neural network.
Since GDFS-DM simply outputs the action $a\in[d]\setminus S$ that would yield largest information gain (and not the actual information gain itself), we can not use the heuristic discussed earlier for the soft-budget case. Instead, we stop when $H(\mathbf{y}\mid x_{S})<\alpha$ (based on $p(\mathbf{y}\mid x_{S})$ output by predictor $f(S,x_{S})$).
@@
DIME-DM (Gadgil et al., 2024). This approach directly estimates the CMI in a discriminative way. Extending the variational estimators of (Chattopadhyay et al., 2023; Covert et al., 2023), introduces a new objective whose optimum equals the true CMI. The practical implementation uses two jointly trained networks: a predictor $f_{\theta}$ and a value network $g_{\phi}$ that outputs $I(\mathbf{y};\mathbf{x}_{i}\mid x_{S})$ for each unobserved feature. When $f_{\theta}$ approaches the Bayes classifier, the expected reduction in cross-entropy achieved by adding the feature $i$ matches its CMI; training $g_{\phi}$ to reproduce this reduction yields a consistent estimator. After training, $g_{\phi}$ selects the next feature by choosing the largest predicted CMI.
D.2. Non-myopic methods via reinforcement learning
As discussed in Section˜2.4, AFA can be formulated as a Markov decision process (MDP) (Dulac-Arnold et al., 2011; Aronsson et al., 2026). Consequently, one can use reinforcement learning (RL) to find non-myopic policies $\pi$ that acquire feature that maximize long-term reward. This approach enables the discovery of policies that can outperform the myopic methods discussed above. Below, we describe all RL-based approaches used in the benchmark.
Unless otherwise stated, all RL-based methods are adapted to the hard and soft-budget settings as follows. All methods are adapted to the hard-budget setting by simply setting $\alpha=0$ in the reward (See Section˜2.4). For the soft-budget setting, we set $\alpha>0$, and set the budget $B=c([d])$ (cost of acquiring all features).
@@ D.2. Non-myopic methods via reinforcement learning
JAFA-MFRL (Shim et al., 2018). This method uses the RNN-based and order-invariant set encoding introduced in (Vinyals et al., 2016) to represent the subsets of acquired features, significantly reducing the complexity of the search space. The approach jointly trains an RL agent and a classifier, ensuring that the feature selection policy is optimized with respect to the classification objective. To solve the underlying MDP, the method uses deep Q-Networks (DQN) (Mnih et al., 2015), enabling efficient policy learning even in high-dimensional feature spaces. The reward function is exactly as in Section˜2.4.
@@
ODIN-MBRL (Zannone et al., 2019). This is a model-based RL framework for AFA. ODIN uses a pre-trained PVAE (Ma et al., 2019) to approximate the conditional distribution of unobserved features $p(\mathbf{x}_{U}\mid x_{S})$ for $U=[d]\setminus S$. Since the MDP transition dynamics can be derived from this distribution, ODIN performs model-based rollouts to simulate feature acquisition trajectories, improving data efficiency and robustness to missing features. In our experiments, we performed rollouts to augment the training set by $200\%$.
@@
In contrast to JAFA-MFRL, this method uses a dense reward function. Let $y_{\text{true}}$ be the true class of instance $x$ and $p_{f}(y_{\text{true}}\mid x_{S})$ be the probability of this class given by the predictor $f(x_{S})$. Then, for any state $(S,x_{S})$ and action $a\in\mathcal{A}_{B}(S)$
(4) $r((S,x_{S}),a)=p_{f}(y_{\text{true}}\mid x_{S},x_{a})-\alpha c(S\cup\{a\})$
which motivates the agent to select the most informative features first. This can alleviate issues with sparse rewards, but may bias the method to more myopic policies (see Section 2.4 for more discussion).
The policy is trained using Proximal Policy Optimization (PPO) (Schulman et al., 2017).
@@
ODIN-MFRL The key strength of ODIN-MBRL is its performance on smaller datasets (Zannone et al., 2019). However, since most of the datasets in this benchmark are relatively large, we decided to also evaluate a variant of ODIN where the agent is trained only on the original dataset, without using generative rollouts.
@@
OL-MFRL+no-mask (Kachuee et al., 2019). This is another model-free and DQN-based method. The main characteristic of it is the reward function. For any state $(S,x_{S})$ and action $a\in\mathcal{A}_{B}(S)$, the reward is defined as
(5) $r((S,x_{S}),a)=\frac{\|\textit{Cert}(x_{S\cup a})-\textit{Cert}(x_{S})\|}{c(a)}$
@@
which does not depend on the true label but only on the classifier predictions. Here Cert is a vector of class probabilities, calculated by averaging over multiple forward passes with dropout. Furthermore, the method uses a coupled PQ-network where the output of the predictor (P-network) is fed into the action-value estimator (Q-network).
This method was originally introduced for the hard-budget setting. Although it effectively adapts to non-uniform costs due to the denominator, every reward is non-negative, providing no incentive for the agent to stop early. Consequently, adapting it to the soft-budget setting would require substantial changes to the reward design and we therefore omit this method from the soft-budget experiments. Our goal is to benchmark existing AFA methods as they are, and avoid major modifications that could reasonably be viewed as defining a different method.
OL-MFRL One major drawback with OL-MFRL+no-mask is that it cannot distinguish between missing features and features with a value of 0. Hence, OL-MFRL is a variation of the original OL-MFRL+no-mask method where we simply concatenate the feature mask to the features in the input.
D.3. Oracle-based methods (non-myopic)
@@
Rather than learning policies through reinforcement learning, oracle-based methods directly approximate an optimal acquisition policy by estimating the underlying data distributions. This approach offers a middle ground between myopic methods and computationally expensive RL training.
@@
AACO (Valancius et al., 2024) formulates AFA as a subset optimization problem. Instead of greedily selecting features one by one, ACO considers acquiring entire subsets of features simultaneously. The optimal policy selects the subset $o^{\prime}\subseteq U$ that minimizes the expected acquisition cost plus the prediction loss,
@@
(6) $u(x_{o},o)=\operatorname*{argmin}_{o^{\prime}\subseteq U}\mathbb{E}_{y,x_{o^{\prime}}\mid x_{o}}\left[\ell\left(\hat{y}(x_{o},x_{o^{\prime}}),y\right)\right]+\alpha|o^{\prime}|,$
@@
In practice, computing this expectation requires knowledge of the joint distribution $p(y,x_{o^{\prime}}\mid x_{o})$, which is unknown. AACO addresses this by using $k$-nearest neighbors density estimation to approximate the required distributions. Furthermore, for computational tractability in high-dimensional spaces, AACO samples a random subset of possible feature combinations rather than evaluating all $2^{|\mathcal{O}|}$ possibilities.
This approach is non-myopic since it jointly optimizes over multiple features rather than selecting them sequentially, yet it avoids the training complexity of RL methods by directly approximating the optimal oracle policy.
The method naturally supports the soft-budget setting by terminating early whenever $u(x_{o},o)=\emptyset$, that is, when stopping and predicting with the currently acquired features achieves the highest utility. In the hard-budget setting, early termination is disallowed, and the method continues selecting features until the budget $B$ is exhausted.
AACO+NN AACO+NN is a neural approximation of AACO. We first generate expert trajectories from a trained AACO oracle. Each trajectory contains states represented by masked features and feature masks, and actions represented by either "stop" or "acquire selection jj". We then train a policy network with behavioral cloning to imitate these oracle actions. At inference time, this policy replaces repeated $k$-NN subset evaluation, while predictions are still made by the same classifier as in AACO.
D.4. Static (non-adaptive) feature selection methods